\begin{thebibliography}{99}
\bibitem{AB12}
M. Ahmia, H. Belbachir, Preserving log-convexity for generalized
Pascal triangles, Electron. J. Combin. 19 (2) (2012), Research Paper
16.
\bibitem{AB10}
M. Ahmia, H. Belbachir, Preserving log-concavity and generalized
triangles, AIP Conf. Proc. 1264, 81 (2010).
\bibitem{Aig98}
M. Aigner, Motzkin numbers, European J. Combin. 19 (1998) 663--675.
\bibitem{Aig991}
M. Aigner, Catalan-like numbers and determinants. J. Combin. Theory,
Ser. A 87 (1999) 33--51.
\bibitem{Aig99}
M. Aigner, A characterization of the Bell numbers, Discrete Math.
205 (1999) 207--210.
\bibitem{Aig01}
M. Aigner, Catalan and other numbers--a recurrent theme, in: H.
Crapo, D. Senato (Eds.), Algebraic Combinatorics and Computer
Science, Springer, Berlin, 2001, pp. 347--390.
\bibitem{Aig08}
M. Aigner, Enumeration via ballot numbers, Discrete Math. 308 (2008)
2544--2563.



%J. Bonin, L. Shapiro, R. Simion, Some $q$-analogues of the
%Schr\"oder numbers arising from combinatorial statistics on lattice
%paths, J. Statist. Plann. Inference 34 (1993) 35--55.
%\bibitem{Bona12}
%M. B\'ona, Combinatorics of Permutations, second edition, CRC Press,
%Boca Raton, FL, 2012.

%\bibitem{Bon02}
%M. B\'ona, Symmetry and Unimodality in Stack sortable permutations,
%J. Combin. Theory Ser. A 98 (2002) 201--209; Corrigendum: J. Combin.
%Theory Ser. A 99 (2002) 191--194.
%\bibitem{BA00}
%M. B\'ona and R. Ehrenborg, A combinatorial proof of the
%log-concavity of the numbers of permutations with $k$ runs, J.
%Combin. Theory Ser. A 90 (2000) 293--303.
%\bibitem{Bon04}
%M. B\'ona, Combinatorics of permutations, Chapman \& Hall/CRC, Boca
%Raton, FL, 2004.
%\bibitem{Bra06}
%P. Br\"and\'en, On linear transformations preserving the P\'olya
%frequency property, Trans. Amer. Math. Soc. 358 (2006) 3697--3716.
%\bibitem{Branden09}
%P. Br\"{a}dr\'{e}n, Iterated sequeces and the geometry of zeros, J.
%Reine Angew. Math, 658 (2011) 115--131.
\bibitem{Bre89}
F. Brenti, Unimodal, log-concave, and P\'olya frequency sequences in
combinatorics,  Mem. Amer. Math. Soc. 413 (1989).
\bibitem{Bre94}
F. Brenti, Log-concave and unimodal sequences in algebra,
combinatorics, and geometry: an update, Contemp. Math. 178 (1994)
71--89.
%\bibitem{Bre94EuJC}
%F. Brenti, $q$-Eulerian polynomials arising from Coxeter groups,
%European J. Combin. 15 (1994) 417--441.
%\bibitem{Bre95}
%F. Brenti, Combinatorics and total positivity, J. Combin. Theory
%Ser. A 71 (1995) 175--218.

%\bibitem{Bre96}
%F. Brenti, The applications of total positivity to combinatorics,
%and conversely, Math. Appl. 359 (1996) 451--473.
%\bibitem{Bre00}
%F. Brenti, A class of $q$-symmetric functions arising from plethysm,
%J. Combin. Theory Ser. A 91 (2000) 137--170.
\bibitem{But90}
L.M. Butler, The q-log concavity of q-binomial coeffcients, J.
Combin. Theory Ser. A 54 (1990) 54--63.
%\bibitem{BF07}
%L.M. Butler, W.P. Flanigan, A note on log-convexity of q-Catalan
%numbers. Ann. Comb. 11 (2007) 369--373.
%\bibitem{Cal00}
%D. Callan,
%Notes on Motzkin and Schr\"oder numbers, available at\\
%\verb|http://www.stat.wisc.edu/~callan/papersother/|
%\bibitem{Car71}
%L. Carlitz, Note on the numbers of Jordan and Ward, Duke Math. 38
%(1971) 783--790.
%\bibitem{CTWY10} W.Y.C. Chen, R.L. Tang, L.X.W. Wang, A.L.B. Yang,
%The q-log-convexity of the Narayana polynomials of type B, Adv. in
%Appl. Math. 44(2) (2010) 85--110.
\bibitem{Com74}
L. Comtet, Advanced combinatorics, D. Reidel Publishing Co.,
Dordrecht, 1974.
\bibitem{CLW15}
X. Chen, H. Liang, Y. Wang, Total positivity of recursive matrices,
Linear Algebra Appl. 471 (2015) 383--393.
%\bibitem{CLW152}
%X. Chen, H. Liang, Y. Wang, Total positivity of Riordan arrays,
%European J. Combin. 46 (2015) 68--74.


\bibitem{CWY10}
W.Y.C. Chen, L.X.W. Wang, A.L.B. Yang, Schur positivity and the
$q$-log-convexity of the Narayana polynomials, J. Algebraic Combin.
32(3) (2010) 303--338.

\bibitem{CTWY10}
W.Y.C. Chen, R.L. Tang, L.X.W. Wang, A.L.B. Yang, The
q-log-convexity of the Narayana polynomials of type B, Adv. in Appl.
Math. 44(2) (2010) 85--110.
\bibitem{CWY11}
W.Y.C. Chen, L.X.W. Wang, A.L.B. Yang, Recurrence relations for
strongly q-log-convex polynomials, Canad. Math. Bull. 54 (2011)
217--229.
%\bibitem{DP49}
%H. Davenport and G. P\'olya, On the product of two power series,
%Canadian J. Math. 1 (1949) 1--5.
\bibitem{CKS12}
G.-S. Cheon, H. Kim and L.W. Shapiro, Combinatorics of Riordan
arrays with iden- tical A and Z sequences, Discrete Math. 312 (2012)
2040--2049.

\bibitem{ES00}
R. Ehrenborg, E. Steingrimsson, The excedance set of a permutation,
Adv. Appl. Math. 24 (2000) 284--299.


%\bibitem{Do73}
%T.A. Dowling, A class of geometric lattices based on ¡¥nite groups,
%J. Combin. Theory Ser. B 14 (1973) 61--86; Erratum, J. Combin.
%Theory Ser. B 15 (1973) 211.
\bibitem{DP49}
H. Davenport and G. P\'olya, On the product of two power series,
Canad. J. Math. 1 (1949) 1--5.

%\bibitem{DV03}
%T. Do\v sli\'c and D. Veljan, Calculus proofs of some combinatorial
%inequalities, Math. Inequal. Appl. 6 (2003) 197--209.
%\bibitem{DSV04}
%T. Do\v sli\'c, D. Svrtan and D. Veljan, Enumerative aspects of
%secondary structures, Discrete Math. 285 (2004) 67--82.

%\bibitem{Dos06}
%T. Do\v sli\'c, Logarithmic behavior of some combinatorial
%sequences, Discrete Math. 308 (2008) 2182--2212.
%\bibitem{Dow73}
%T. A. Dowling, A class of geometric lattices based on finite groups,
%J. Combin. Theory Ser. B 14 (1973) 61--86; Erratum, J. Combin.
%Theory Ser. B 15 (1973) 211.
%\bibitem{E94}
%K. Engel, On the average rank of an element in a filter of the
%partition lattice, J. Combin. Theory Ser. A 65 (1994) 67--78.

%\bibitem{Fisk08}
%S. Fisk, Polynomials, roots, and interlacing, arXiv:math/0612833.
%\bibitem{FS70}
%D. Foata, M. Sch\"utzenberger, Th\'eorie G\'eom\'etrique des
%Polyn\^omes Euleriens, Lecture Notes in Mathematics, No. 138,
%Springer-Verlag, Berlin/New York, 1970.
%\bibitem{G59}
%F.R. Gantmacher, The Theory of Matrices, vol. II, Chelsea Publ., New
%York, 1959.

%\bibitem{HLP52}
%G.H. Hardy, J.E. Littlewood, G. P\'olya, Inequalities, Cambridge
%University Press, Cambridge, 1952.

%\bibitem{Har67}
%L.H. Harper, Stirling behavior is asymptotically normal, Ann. Math.
%Statist. 38 (1967) 410--414.
%\bibitem{H13}
%T.-X. He, Parametric Catalan numbers and Catalan triangles, Linear
%Algebra Appl. 438 (2013) 1467--1484.

%\bibitem{HS09}
%T.-X. He and R. Sprugnoli, Sequence characterization of Riordan
%arrays, Discrete Math. 309 (2009) 3962--3974.
\bibitem{Kar68}
S. Karlin, Total Positivity, Vol. I, Stanford University Press,
Stanford, 1968.
\bibitem{Kra}
C. Krattenthaler, On the q-log-concavity of Gaussian binomial
coefficients, Monatsh. Math. 107 (1989), 333--339.
%\bibitem{Kur72}
%D.C. Kurtz, A note on concavity properties of triangular arrays of
%numbers, J. Combin. Theory Ser. A 13 (1972) 135--139.
%\bibitem{Kur73}
%D.E. Knuth, The Art of Computer Programming, vol. 3, Fundamental
%Algorithms, Addison-Wesley, Reading, MA, 1973.

\bibitem{Le90}
P. Leroux, Reduced matrices and q-log-concavity properties of
q-Stirling numbers, J. Combin. Theory Ser. A 54 (1990) 64--84.
%\bibitem{LW-RZP}
%L.L. Liu, Y. Wang, A unified approach to polynomial sequences with
%only real zeros, Adv. in Appl. Math. 38(4) (2007) 542--560.

%\bibitem{LZ14}
%Z. Lin, J. Zeng, Positivity properties of Jacobi-Stirling numbers
%and generalized Ramanujan polynomials, Adv. in Appl. Math. 53
%(2014), 12--27.
\bibitem{LW07}
L.L. Liu, Y. Wang, On the log-convexity of combinatorial sequences,
Adv. in. Appl. Math. 39 (2007) 453--476.
%\bibitem{LW-RZP}
%.L. Liu, Y. Wang, A unified approach to polynomial sequences with
%only real zeros, Adv. in Appl. Math. 35 (2007) 542--560.
%\bibitem{LZ15}
%L.L. Liu, B.-X. Zhu, Strong $q$-log-convexity of the Eulerian
%polynomials of Coxeter groups, Discrete Math. 338 (2015) 2332--2340.
%\bibitem{PS64}
%G. P\'{o}lya, G. Szeg\"o, Problems and Theorems in Analysis, Vol.
%II, 3rd ed., 1964.
%\bibitem{Ma13}
%S.M. Ma, Enumeration of permutations by number of alternating runs,
%Discrete Math. 313 (2013) 1816--1822.
%\bibitem{Ma121}
%S.M. Ma, Derivative polynomials and enumeration of permutations by
%number of interior and left peaks, Discrete Math. 312 (2012)
%405--412.

\bibitem{MRSV97}
D. Merlini, D.G. Rogers, R. Sprugnoli, M.C. Verri, On some
alternative characterizations of Riordan arrays, Canad. J. Math.
49(2)(1997)301--320.
%\bibitem{MW08}
%S.M. Ma, Y. Wang, q-Eulerian polynomials and polynomials with only
%real zeros. Electron. J. Combin. 15 (2008), no. 1, Research Paper
%17.

%\bibitem{Monge12}
%P. Mongelli, Total positivity properties of Jacobi--Stirling
%numbers, Adv. in. Appl. Math. 48 (2012) 354--364.
%\bibitem{Na62}
%T.V. Narayana, An analogue of the multinomial theorem. Canad. Math.
%Bull. 5 (1962) 43--50.
%\bibitem{Pit97}
%J. Pitman, Probabilistic bounds on the coefficients of polynomials
%with only real zeroes, J. Combin. Theory Ser. A 77 (1997) 279--303.
%\bibitem{Pos81}
%M.M. Postnikov, Stable Polynomials, Nauka, Moscow, 1981 (in
%Russian).
%\bibitem{Sag88}
%B.E. Sagan, Inductive and injective proofs of log concavity results,
%Discrete Math. 68 (1988) 281--292.
%\bibitem{Ri}
%J. Riordan, Combinatorial Identities, New York, 1979.
%\bibitem{R78}
%D.G. Rogers, Pascal triangles, Catalan numbers and renewal arrays,
%Discrete Math. 22 (1978) 301--310.


\bibitem{Sl}
N.J.A. Sloane, The On-Line Encyclopedia of Integer Sequences,
http://oeis.org.

%\bibitem{Sag98}
%B. E. Sagan, Unimodality and the reflection principle, Ars Combin.
%48 (1998) 65--72.
\bibitem{Sag92}
B.E. Sagan, Inductive proofs of $q$-log concavity, Discrete Math. 99
(1992) 289--306.
\bibitem{Sag921}
B.E. Sagan, Log concave sequences of symmetric functions and analogs
of the Jacobi-Trudi determinants, Trans. Amer. Math. Soc. 329 (1992)
795--811.


%\bibitem{Sta08}
%R.P. Stanley, Longest alternating subsequences of permutations,
%Michigan Math. J. 57 (2008) 675--687.
\bibitem{Sta89}
R.P. Stanley, Log-concave and unimodal sequences in algebra,
combinatorics, and geometry, Ann. New York Acad. Sci. 576 (1989)
500--534.

%\bibitem{Shap03}
%L.W. Shapiro, Bijections and the Riordan group, Theoret. Comput.
%Sci. 307(2) (2003) 403--413.
%\bibitem{SGWW91}
%L.W. Shapiro, S. Getu, W.-J. Woan, L.C. Woodson, The Riordan group,
%Discrete Appl. Math. 34(1--3) (1991) 229--239.
\bibitem{Sha76}
L.W. Shapiro, A Catalan triangle, Discrete Math. 14 (1976) 83--90.
%\bibitem{Spr94}
%R. Sprugnoli, Riordan arrays and combinatorial sums, Discrete Math.
%132 (1994) 267--290.

%\bibitem{Spr11}
%R. Sprugnoli, Combinatorial sums through Riordan arrays, J. Geom.
%101 (2011) 195--210.
%\bibitem{Tan75}
%S. Tanny, On some numbers related to the Bell numbers, Canad. Math.
%Bull. 17 (1975) 733--738.
%\bibitem{SW08}
%X.-T. Su, Y. Wang, On the unimodality problems in Pascal triangle,
%Electron. J. Combin. 15 (2008), Research Paper 113.
\bibitem{SWY11}
X.T. Su, Y. Wang, Y.N. Yeh, Unimodality Problems of Multinomial
Coefficients and Symmetric Functions, Electron. J. Combin. 18(1)
(2011), Research Paper 73.
%\bibitem{S13}
%Z.-W. Sun, Congruences for Franel numbers, Adv. in. Appl. Math.
%51(4) (2013) 524--535.
%\bibitem{S87}
%L.A. Sz\'{e}kely, Holiday numbers: sequences resembling to the
%Stirling numbers of the second kind, Acta Sci. Math. 48 (1985)
%459--467.
%\bibitem{Wag91}
%D.G. Wagner, The partition polynomials of a finite set system, J.
%Combin. Theory Ser. A 56 (1991) 138--159.
%\bibitem{Wag92}
%D.G. Wagner, Total positivity of Hadamard products, J. Math. Anal.
%Appl. 163 (1992) 459--483.
%\bibitem{Wag92}
%D.G. Wagner, Total positivity of Hadamard products, J. Math. Anal.
%Appl. 163 (1992) 459--483.
\bibitem{W03}
Y. Wang, Linear transformations preserving log-concavity, Linear
Algebra and its Applications 359 (2003) 161--167.

\bibitem{WY07}
Y. Wang and Y.-N. Yeh, Log-concavity and LC-positivity, J. Combin.
Theory Ser. A 114 (2007) 195--210.

%\bibitem{Chen081}
%W.Y.C. Chen, Log-concavity and q-Log-convexity Conjectures on the
%Longest Increasing Subsequences of Permutations,  arXiv:
%0806.3392v2.
%\bibitem{WZ14}
%Y. Wang, Z.-H. Zhang, Log-convexity of Aigner-Catalan-Riordan
%numbers, Linear Algebra Appl. 463 (2014) 45--55.


\bibitem{Zhu13}
B.-X. Zhu, Log-convexity and strong $q$-log-convexity for some
triangular arrays, Adv. in. Appl. Math. 50(4) (2013) 595--606.
\bibitem{Zhu14}
B.-X. Zhu, Some positivities in certain triangular array, Proc.
Amer. Math. Soc. 142(9) (2014) 2943--2952.
%\bibitem{Zhu16}
%B.-X. Zhu, Log-concavity and strong $q$-log-convexity for  Riordan
%arrays and recursive matrices, Proceedings of the Royal Society of
%Edinburgh, Section A: Mathematics, to appear, 2016.
\bibitem{ZS15}
B.-X. Zhu, H. Sun, Linear transformations preserving the strong
$q$-log-convexity of polynomials,  Electron. J. Combin., 22(3)
(2015)  Paper 3.27.
\end{thebibliography}

\documentclass[12pt,reqno]{article}
\usepackage{amsfonts}
\usepackage{mathrsfs}

\usepackage{amsmath,amsthm,amsfonts,amssymb,latexsym,mathrsfs,color}

\usepackage[colorlinks=true,
linkcolor=webblue, filecolor=webbrown, citecolor=webred ]{hyperref}
\definecolor{webblue}{rgb}{0,.5,0}
\definecolor{webred}{rgb}{0,.5,0}
\definecolor{webbrown}{rgb}{.6,0,0}

\newtheorem{thm}{Theorem}[section]

\newtheorem{cor}[thm]{Corollary}
\newtheorem{prop}[thm]{Proposition}

\newtheorem{ex}[thm]{Example}

\newtheorem{f}{Fact}[section]

\theoremstyle{definition}

\theoremstyle{remark}
\newtheorem{rem}[thm]{Remark}

\numberwithin{equation}{section}

\title{Linear transformations and strong $q$-log-concavity for certain combinatorial triangle
\thanks{Supported partially by the National Natural Science Foundation of China (Grant No. 11571150).
\newline\hspace*{5mm}
   {\it Email address:}\quad
    bxzhu@jsnu.edu.cn (B.-X. Zhu)}}
\author{Bao-Xuan Zhu}
\date{\footnotesize School of Mathematics and Statistics,
         Xuzhou Normal University,
         Xuzhou 221116, PR China}

\begin{document}

\maketitle

\begin{abstract}
It is well-known that the binomial transformation  preserves the
log-concavity property and log-convexity property. Let
$\binom{a+n}{b+k}$ be the binomial coefficients and $\binom{n,k}{j}$
be defined by
$(b_0+b_1x+\cdots+b_kx^{k})^n:=\sum_{j=0}^{kn}\binom{n,k}{j}x^j,$
where the sequence $(b_i)_{0\leq i\leq k}$ is log-concave. In this
paper, we prove that the linear transformation
$$y_n(q)=\sum_{k=0}^n\binom{a+n}{b+k}x_k(q)$$ preserves the strong $q$-log-concavity
property for any fixed nonnegative integers $a$ and $b$, which
strengthens and gives a simple proof of results of Ehrenborg and
Steingrimsson, and Wang, respectively, on linear transformations
preserving the log-concavity property. We also show that the linear
transformation
$$y_n=\sum_{i=0}^{kn}\binom{n,k}{j}x_i$$ not only preserves the log-concavity property, but also preserves the log-convexity
property, which extends the results of Ahmia and Belbachir about the
$s$-triangle transformation preserving the log-convexity property
and log-concavity property. Let $[A_{n,k}(q)]_{n, k\geq0}$ be an
infinite lower triangular array of polynomials in $q$ with
nonnegative coefficients satisfying the recurrence
\begin{eqnarray*}\label{re}
A_{n,k}(q)=f_{n,k}(q)\,A_{n-1,k-1}(q)+g_{n,k}(q)\,A_{n-1,k}(q)+h_{n,k}(q)\,A_{n-1,k+1}(q),
\end{eqnarray*}
for $n\geq 1$ and $k\geq 0$, where $A_{0,0}(q)=1$,
$A_{0,k}(q)=A_{0,-1}(q)=0$ for $k>0$. We present criterions for the
strong $q$-log-concavity of the sequences in each row of
$[A_{n,k}(q)]_{n, k\geq0}$. As applications, we get the strong
$q$-log-concavity or the log-concavity of the sequences in each row
of many well-known triangular arrays, such as the Bell polynomials
triangle, the Eulerian polynomials triangle and the Narayana
polynomials triangle in a unified approach.
\bigskip\\
{\sl MSC:}\quad 05A10; 05A20; 11B73; 15B36.
\bigskip\\
{\sl Keywords:}\quad Combinatorial triangles; Log-concavity; Strong
$q$-log-concavity; Linear transformations
\end{abstract}

\section{Introduction}
\hspace*{\parindent}
Let $(a_n)_{n\geq0}$ be a sequence of nonnegative real numbers. It
has no {\it internal zeros }, {\it i.e.}, $a_j\neq0$ if
$a_ia_k\neq0$ for $i< j< k$. Thoughout this paper, all sequences
have no internal zeros. We call it {\it log-concave} if
$a_{k-1}a_{k+1}\le a_k^2$ for all $k\ge 1$. The log-concave
sequences arise often in combinatorics, algebra, geometry, analysis,
probability and statistics and have been extensively investigated,
see Stanley~\cite{Sta89} and Brenti~\cite{Bre94} for details.

For two polynomials with real coefficients $f(q)$ and $g(q)$, denote
$f(q)\geq_q g(q)$ if the difference
$f\left(q\right)-g\left(q\right)$ has only nonnegative coefficients.
For a polynomial sequence $(f_n(q))_{n\geq 0}$, it is called {\it
$q$-log-concave} suggested by Stanley if
$$f_n(q)^2\geq_q f_{n+1}(q)f_{n-1}(q),$$ for $n\geq 1$ and is called {\it strongly $q$-log-concave} introduced by Sagan if
$$f_{n+1}(q)f_{m-1}(q)\geq_q f_n(q)f_m(q),$$
for any $m\geq n\geq1$. Obviously, the strong $q$-log-concavity
implies the $q$-log-concavity.  The q-log-concavity of polynomials
have been extensively studied, see Butler \cite{But90},
Krattenthaler \cite{Kra}, Leroux \cite{Le90}, Sagan
\cite{Sag92,Sag921}, and Su, Wang and Yeh \cite{SWY11} for instance.
If we reverse the inequalities in above definitions, then we have
the concepts for log-convexity, $q$-log-convexity and strong
$q$-log-convexity, respectively. Reader can refer to Chen {\it et
al.} \cite{CWY10,CTWY10,CWY11}, Liu and Wang \cite{LW07}, Zhu
\cite{Zhu13,Zhu14}, and Zhu and Sun \cite{ZS15} for the strong
$q$-log-convexity.

The linear transformations are often used to study the log-concavity
and log-convexity. For instance, it was proved that the binomial
transformation $b_n=\sum_{k=0}^n\binom{n}{k}a_k$ for $n\geq0$
preserves the log-concavity property, {\it i.e.}, log-concavity of
$(a_n)_{n\geq 0}$ implies that of $(b_n)_{n\geq 0}$ (see
\cite[Theorem 2.5.7]{Bre89} or \cite[Theorem 7.3]{Kar68} for
instance) and log-convexity property \cite{LW07}. More generally,
the log-convexity property and log-concavity property are preserved
under the binomial convolution, see Davenport and
P\'olya~\cite{DP49} and Wang and Yeh~\cite{WY07}, respectively.
Using certain positivity method, Ehrenborg and Steingrimsson
\cite{ES00} also showed that the linear transformation
$$y_n=\sum_{k=0}^n\binom{n+1}{k}x_k$$ preserves the log-concavity
property. Furthermore, using the algebraical method, Wang \cite{W03}
proved one more general result that the linear transformation
$$y_n=\sum_{k=0}^n\binom{a+n}{b+k}x_k$$ preserves the log-concavity
property for fixed nonnegative integers $a$ and $b$.  Motivated by
these, we prove the following stronger result.
\begin{thm}\label{thm+binom}
Let $a$ and $b$ be two
nonnegative integers. Then the linear transformation
$$y_n(q)=\sum_{k=0}^n\binom{a+n}{b+k}x_k(q)$$ preserves the strong $q$-log-concavity
property. In particular, it preserves the log-concavity property of
sequences.
\end{thm}
The $s$-triangle $\binom{n}{j}_s$ is a generalization of the Pascal
triangle, which is given by the ordinary multinomial coefficients
\cite{Com74}:
$$(1+x+\cdots+x^{s})^n:=\sum_{j=0}^{sn}\binom{n}{j}_sx^j.$$
In \cite{AB12,AB10}, Ahmia and Belbachir also demonstrated that the
log-convexity property and log-concavity property are preserved
under the $s$-triangle transformation. Motivated by this, we will
consider a more general triangle in the following. Let $k$ be a
positive integer. Assume that the nonnegative sequence $(b_i)_{0\leq
i\leq k}$ is log-concave. Define a more generalized triangle
$\binom{n,k}{j}$:
$$(b_0+b_1x+\cdots+b_kx^{k})^n:=\sum_{j=0}^{kn}\binom{n,k}{j}x^j.$$
It is obvious that the triangle $[\binom{n,k}{j}]_{n,j\geq0}$ has
the following recurrence relations:\begin{eqnarray}
\binom{n,k}{j}&=&\sum_{i=0}^k\binom{n-1,k}{j-i}b_i,\label{recur+1}\\
\binom{n,k}{j}&=&\sum_{i=0}^{sk}\binom{n-s,k}{i}\binom{s,k}{j-i}.\label{recur+s}
\end{eqnarray}
In fact, the triangle $[\binom{n,k}{j}]_{n,j\geq0}$ generalizes many
famous triangles. For instance,
\begin{itemize}
\item[\rm (i)]
If both $k=b_0=b_1=1$, then the triangle
$[\binom{n,k}{j}]_{n,j\geq0}$ turns out be the Pascal triangle.
\item[\rm (ii)]
If $b_0=b_1=\ldots=b_k=1$, then the triangle
$[\binom{n,k}{j}]_{n,j\geq0}$ turns out be the s-triangles given by
the ordinary multinomials, see \cite[A027907 for s = 2, A008287 for
s = 3 and A035343 for s = 4]{Sl}.
\end{itemize}
We also prove the next stronger result in a unified approach.
\begin{thm}\label{thm+three}
Let $k$ be any fixed positive integer and $\binom{n,k}{j}$ be as
above. Then the linear transformation
$$y_n=\sum_{i=0}^{kn}\binom{n,k}{j}x_i$$ not only preserves the log-concavity property, but also preserves the log-convexity property.
\end{thm}

In \cite{Zhu13}, Zhu defined a triangular array as follows. Let
$[A_{n,k}(q)]_{n, k\geq 0}$ be an infinite lower triangular array
defined by the recurrence
\begin{eqnarray}\label{rre}
A_{n,k}(q)=f_{k}(q)\,A_{n-1,k-1}(q)+g_{k}(q)\,A_{n-1,k}(q)+h_{k}(q)\,A_{n-1,k+1}(q)
\end{eqnarray}
for $n\geq 1$ and $k\geq 0$, where $A_{0,0}(q)=1$,
$A_{0,k}(q)=A_{0,-1}(q)=0$ for $k>0$. This triangular array consists
of polynomials in $q$ and it will turn out to be the array with
entries be real numbers for any fixed $q\geq0$. In fact, this
triangular array (\ref{rre}) unifies many well-known combinatorial
triangles. The following are some basic examples.

\begin{ex}\label{exm}
(1) The Catalan triangle of Aigner \cite{Aig991} is
$$C=[C_{n,k}]_{n,k\geq0}=\left[
      \begin{array}{rrrrr}
        1 &  &  &  &  \\
        1 & 1 &   &   &\\
        2 & 3 & 1 &   & \\
        5 & 9 & 5 & 1 &   \\
       \vdots &  &  &  & \ddots \\
      \end{array}
    \right],$$
where $C_{n+1,k}=C_{n,k-1}+2\,C_{n,k}+C_{n,k+1}$ and
$C_{n+1,0}=C_{n,0}+C_{n,1}$. The numbers in the $0$th column are the Catalan numbers $C_n$. \\
(2) The Catalan triangle of Shaprio \cite{Sha76} is
$$C'=[C'_{n,k}]_{n,k\geq0}=\left[
      \begin{array}{rrrrr}
        1 &  &  &  &  \\
        2 & 1 &   &   &\\
        5 & 4 & 1 &   & \\
        14 & 14 & 6 & 1 &   \\
       \vdots &  &  &  & \ddots \\
      \end{array}
    \right],$$
where $C'_{n+1,k}=C'_{n,k-1}+2\,C'_{n,k}+C'_{n,k+1}$ for $k\geq0$. The numbers in the $0$th column are the Catalan numbers $C_n$. \\
(3) The Motzkin triangle \cite{Aig98,Aig991} is
  $$M=[M_{n,k}]_{n,k\geq0}=\left[
  \begin{array}{rrrrr}
                   1 &  &  &  &  \\
                   1 & 1 &  &  &  \\
                   2 & 2 & 1 &  &  \\
                   4 & 5 & 3 & 1 &  \\
                   \vdots &  &  &  & \ddots \\
                 \end{array}
  \right],$$
where $M_{n+1,k}=M_{n,k-1}+M_{n,k}+M_{n,k+1}$ and
$M_{n+1,0}=M_{n,0}+M_{n,1}$. The numbers in the $0$th column are the Motzkin numbers $M_n$.\\
(4) The large Schr\"{o}der triangle \cite{CKS12} is
 $$s=[s_{n,k}]_{n,k\geq0}=\left[
 \begin{array}{rrrrr}
 1 &  &  &  &  \\
 2 & 1 &  &  &  \\
 6 & 4 & 1 &  &  \\
 22 & 16 & 6 & 1 &  \\
 \vdots &  &  &  & \ddots \\
 \end{array}
 \right],$$ where $s_{n+1,k}=s_{n,k-1}+2\,s_{n,k}+2s_{n,k+1}$ and
$s_{n+1,0}=s_{n,0}+2s_{n,1}$. The numbers in the $0$th column are the large Schr\"{o}der numbers $S_n$. \\
(5) The Bell triangle \cite{Aig99} is
  $$B=[B_{n,k}]_{n,k\geq0}=\left[
  \begin{array}{lllll}
                   1 &  &  &  &  \\
                   1 & 1 &  &  &  \\
                   2 & 3 & 1 &  &  \\
                   5 & 10 & 6 & 1 &  \\
                   \vdots &  &  &  & \ddots \\
                 \end{array}
  \right],$$
  where $B_{n+1,k}=B_{n,k-1}+(1+k)\,B_{n,k}+(1+k)\,B_{n,k+1}$ and
$B_{n+1,0}=B_{n,0}+B_{n,1}$. The numbers in the $0$th column are the
Bell numbers.\\
(6) The Bell polynomials triangle $B=[B_{n,k}]_{n,k\geq0}$
\cite{Aig01} is
  $$\left[
  \begin{array}{lllll}
                   1 &  &  &  &  \\
                   q & 1 &  &  &  \\
                   q^2+q & 2q+1 & 1 &  &  \\
                   q^3+3q^2+q & 3q^2+6q+1 & 3q+3 & 1 &  \\
                   \vdots &  &  &  & \ddots \\
                 \end{array}
  \right],$$
  where $B_{n+1,k}=B_{n,k-1}+(q+k)\,B_{n,k}+q(1+k)\,B_{n,k+1}$. The numbers in the $0$th column are the Bell polynomials. In particular, it reduces to the Bell triangle
  \cite{Aig99}
  for $q=1$.\\
(7) The Eulerian polynomials triangle $E=[E_{n,k}]_{n,k\geq0}$
\cite{Aig01} is
  $$\left[
  \begin{array}{lllll}
                   1 &  &  &  &  \\
                   1 & 1 &  &  &  \\
                   q+1 & q+3 & 1 &  &  \\
                   q^2+4q+1 & q^2+10q+7 & 3q+6 & 1 &  \\
                    \vdots &  &  &  & \ddots \\
                 \end{array}
  \right],$$
  where $E_{n+1,k}=E_{n,k-1}+(kq+k+1)\,E_{n,k}+(k+1)^2q\,E_{n,k+1}$. The numbers in the $0$th column are
the Eulerian polynomials.\\
(8) The Narayana polynomials triangle $N=[N_{n,k}]_{n,k\geq0}$
\cite{Aig01} is
  $$\left[
  \begin{array}{lllll}
                   1 &  &  &  &  \\
                   q & 1 &  &  &  \\
                   q^2+q & 2q+1 & 1 &  &  \\
                   q^3+3q^2+q & 3q^2+5q+1 & 3q+2 & 1 &  \\
                    \vdots &  &  &  & \ddots \\
                 \end{array}
  \right],$$
  where $E_{n+1,k}=E_{n,k-1}+(q+1)\,E_{n,k}+q\,E_{n,k+1}$ and $E_{n+1,0}=q E_{n,0}+ q E_{n,1}$. The numbers in the $0$th column are
the Narayana polynomials \cite{Aig01}.
\end{ex}

If $f_k\equiv1$ for $k\geq0$, then it can be reduced to the
recursive matrix and $A_{n,0}(q)$ are called the Catalan-like
numbers, see Aigner~\cite{Aig991,Aig01}. Many combinatorial and
algebraic properties of the triangular array $[A_{n,k}(q)]_{n, k\geq
0}$ have been found. For instance, Aigner
\cite{Aig98,Aig991,Aig01,Aig08} researched various combinatorial
properties of recursive matrices and Hankel matrices of the
Catalan-like numbers. Chen, Liang and Wang \cite{CLW15} considered
the total positivity of recursive matrices. Zhu \cite{Zhu13} gave a
criterion for the strong $q$-log-convexity of the first column of
$[A_{n,k}(q)]_{n,k\geq0}$. However, there is no result for the
strong $q$-log-concavity of the sequence in each row of
$[A_{n,k}(q)]_{n,k\geq0}$. This is our another motivation. In \S4,
we will present some criterions for the strong $q$-log-concavity of
the sequence in each row of $[A_{n,k}(q)]_{n, k\geq0}$.

The remainder part of this paper is arranged as follows. In \S2 and
\S3, we give the proofs of Theorem \ref{thm+binom} and
\ref{thm+three}, respectively. In \S4 we present sufficient
conditions for the strong $q$-log-concavity of rows of the certain
triangle. As applications, we show the log-concavity and the strong
$q$-log-concavity of sequences of rows in many well-known triangular
arrays, such as the Catalan triangles of Aigner and Shaprio, the
large Schr\"oder triangle, the Motzkin triangle, the Bell
polynomials triangle, the Eulerian polynomials triangle and the
Narayana polynomials triangle, and so on, in a unified approach.

\section{Proof of Theorem \ref{thm+binom}}
\begin{proof}
Since a nonnegative sequence $(a_n)_{n\geq0}$ with no internal zeros
is log-concave if and only if $a_ia_j\geq a_{i+1}a_{j-1}$ for $i\geq
j$, the sequence sequence $(a_n)_{n\geq0}$ is strongly
$q$-log-concave. Thus, the second part immediately follows from the
first part for $q=0$. So we only need to show the first part, {\it
i.e.} the linear transformation
$$y_n(q)=\sum_{k=0}^n\binom{a+n}{b+k}x_k(q)$$ preserves the strong $q$-log-concavity
property. We first need the next fact.
\begin{f}\label{sum}
If $(x_n(q))_{n\geq0}$ is strongly $q$-log-concave, then so is
$(x_n(q)+x_{n+1}(q))_{n\geq0}$.
\end{f}

\begin{proof}
Since the sequence $(x_n(q))_{n\geq0}$ is strongly $q$-log-concave,
by the definition, we have $x_i(q)x_j(q)\geq_q x_{i-1}(q)x_{j+1}(q)$
for any $j\geq i\geq0$, which implies that
\begin{eqnarray*}
&&\left[x_{j-1}(q)+x_{j-2}(q)\right]\left[x_{i+1}(q)+x_{i}(q)\right]-\left[x_j(q)+x_{j-1}(q)\right]\left[x_i(q)+x_{i-1}(q)\right]\\
&=&\left[x_{j-1}(q)x_{i+1}(q)-x_j(q)x_i(q)\right]+\left[x_{j-2}(q)x_{i}(q)-x_{j-1}(q)x_{i-1}(q)\right]\\
&&+\left[x_{j-2}(q)x_{i+1}(q)-x_{j}(q)x_{i-1}(q)\right]\\
&\leq_q&0
\end{eqnarray*}
for any $i\geq j\geq0,$ as desired. This proves the fact.
\end{proof}
In the following, we will prove the desired result by induction on
$n$.

If $0\leq n \leq 3$, then we have
\begin{eqnarray}
y_0 (q)&=&x_0(q)\binom{a}{b},\label{0}\\
y_1(q)&=& x_0(q)\binom{a+1}{b} + x_1(q)\binom{a+1}{b+1},\label{1}\\
y_2(q)& =& x_0(q) \binom{a+2}{b}+ x_1(q)\binom{a+2}{b+1}+
x_2(q)\binom{a+2}{b+2},\label{2}\\
y_3(q)& =& x_0(q) \binom{a+3}{b}+ x_1(q)\binom{a+3}{b+1}+
x_2(q)\binom{a+3}{b+2}+x_3(q)\binom{a+3}{b+3},\label{3}
\end{eqnarray}
since $y_n(q)=\sum_{k=0}^n\binom{n+a}{k+b}x_k(q)$. Hence, by
(\ref{0})-(\ref{2}) we obtain that
\begin{eqnarray}\label{eq+}
y_1^2(q)-y_2(q)y_0(q)
&=&x_0^2(q)\left[\binom{a+1}{b}^2-\binom{a}{b}\binom{a+2}{b}\right]\nonumber\\
&&+\left[x_1^2(q)\binom{a+1}{b+1}^2-x_0(q)x_2(q)\binom{a}{b}\binom{a+2}{b+2}\right]\nonumber\\
&&+x_0(q)x_1(q)\left[2\binom{a+1}{b}\binom{a+1}{b+1}-\binom{a}{b}\binom{a+2}{b+1}\right].
\end{eqnarray}
Thus we deduce for $a<b$ that $$y_1^2(q)-y_2(q)y_0(q)\geq_q 0.$$ On
the other hand, by (\ref{eq+}) we also have
$$y_1^2(q)-y_2(q)y_0(q)\geq_q0$$ for $a\geq b$ since
\begin{eqnarray*}
&&\binom{a+1}{b}^2-\binom{a}{b}\binom{a+2}{b}=\binom{a+1}{b-1}\binom{a}{b},\\
&&x_1^2(q)\binom{a+1}{b+1}^2-x_0(q)x_2(q)\binom{a}{b}\binom{a+2}{b+2}\\
&&\geq_q
x_1^2(q)\left[\binom{a+1}{b+1}^2-\binom{a}{b}\binom{a+2}{b+2}\right]=\frac{x_1^2(q)}{b+1}\binom{a+1}{b+2}\binom{a}{b},\\
&&2\binom{a+1}{b}\binom{a+1}{b+1}-\binom{a}{b}\binom{a+2}{b+1}=\frac{a}{a-b+1}\binom{a+1}{b+1}\binom{a}{b}.
\end{eqnarray*}
In the following, we will prove $y_1(q)y_2(q)-y_3(q)y_0(q)\geq_q 0$.
Note for $a\leq b$ that $y_0(q)=0$. So it suffices to consider the
remaining case $a\geq b$. It follows from (\ref{0})-(\ref{3}) we
have
\begin{eqnarray*}
&&y_1(q)y_2(q)-y_3(q)y_0(q)\nonumber\\
&=&\left[x_0(q)\binom{a+1}{b} +
x_1(q)\binom{a+1}{b+1}\right]\left[x_0(q) \binom{a+2}{b}+
x_1(q)\binom{a+2}{b+1}+
x_2(q)\binom{a+2}{b+2}\right]\nonumber\\
&&-x_0(q)\binom{a}{b}\left[ x_0(q) \binom{a+3}{b}+
x_1(q)\binom{a+3}{b+1}+
x_2(q)\binom{a+3}{b+2}+x_3(q)\binom{a+3}{b+3}\right]\nonumber\\
&=&x^2_0(q)\left[\binom{a+1}{b}\binom{a+2}{b}-\binom{a}{b}\binom{a+3}{b}\right]+x_1^2(q)\binom{a+1}{b+1}\binom{a+2}{b+1}\nonumber\\
&&+x_0(q)x_1(q)\left[\binom{a+1}{b}\binom{a+2}{b+1}+\binom{a+1}{b+1}\binom{a+2}{b}-\binom{a}{b}\binom{a+3}{b+1}\right]\nonumber\\
&&+x_0(q)x_2(q)\left[\binom{a+1}{b}\binom{a+2}{b+2}-\binom{a}{b}\binom{a+3}{b+2}\right]\nonumber\\
&&+x_1(q)x_2(q)\binom{a+1}{b+1}\binom{a+2}{b+2}-x_0(q)x_3(q)\binom{a}{b}\binom{a+3}{b+3}\nonumber\\
&\geq_q&x^2_0(q)\left[\binom{a+1}{b}\binom{a+2}{b}-\binom{a}{b}\binom{a+3}{b}\right]\nonumber\\
&&+x_0(q)x_1(q)\left[\binom{a+1}{b}\binom{a+2}{b+1}+\binom{a+1}{b+1}\binom{a+2}{b}-\binom{a}{b}\binom{a+3}{b+1}\right]\nonumber\\
&&+x_0(q)x_2(q)\left[\binom{a+1}{b+1}\binom{a+2}{b+1}+\binom{a+1}{b}\binom{a+2}{b+2}-\binom{a}{b}\binom{a+3}{b+2}\right]\nonumber\\
&&+x_1(q)x_2(q)\left[\binom{a+1}{b+1}\binom{a+2}{b+2}-\binom{a}{b}\binom{a+3}{b+3}\right]\\
&\geq_q&0
\end{eqnarray*}
because $(x_n(q))_{n\geq0}$ is strongly $q$-log-concave and the
following equalities
\begin{eqnarray*}
&&\binom{a+1}{b}\binom{a+2}{b}-\binom{a}{b}\binom{a+3}{b}=\frac{2b}{(a+1-b)(a+3-b)}\binom{a}{b}\binom{a+2}{b},\\
&&\binom{a+1}{b}\binom{a+2}{b+1}+\binom{a+1}{b+1}\binom{a+2}{b}-\binom{a}{b}\binom{a+3}{b+1}=\frac{a(a-b)+a+b}{(a+1-b)(a+2-b)}\binom{a}{b}\binom{a+2}{b+1},\\
&&\binom{a+1}{b+1}\binom{a+2}{b+1}+\binom{a+1}{b}\binom{a+2}{b+2}-\binom{a}{b}\binom{a+3}{b+2}=\frac{ab+2a-b}{(1+b)(2+b)}\binom{a}{b}\binom{a+2}{b+1},\\
&&\binom{a+1}{b+1}\binom{a+2}{b+2}-\binom{a}{b}\binom{a+3}{b+3}=\frac{2(a-b)}{(b+1)(b+3)}\binom{a}{b}\binom{a+2}{b+2}.
\end{eqnarray*}
 Thus, we
obtain that $y_0(q),y_1(q),y_2(q),y_3(q)$ is strongly
$q$-log-concave. So we proceed to the inductive step ($n\geq 4$).

Note that
\begin{eqnarray*}
y_n(q)&=&\sum_{k=0}^n\binom{a+n}{b+k}x_k(q)\\
&=&\sum_{k=0}^{n-1}\binom{a+n-1}{b+k}[x_k(q)+x_{k+1}(q)].
\end{eqnarray*}
Thus by the induction hypothesis and the strong $q$-log-concavity of
$(x_k(q)+x_{k+1}(q))_{k\ge 0}$ by Fact \ref{sum}, we have
$y_0(q),y_1(q),y_2(q),\ldots, y_n(q)$ is strongly $q$-log-concave.
This completes the proof.
\end{proof}
\begin{rem}
If we only consider the log-concavity, we don't need to prove
$y_1y_2-y_0y_3\geq0$ in induction base. Thus, it is obvious that our
proof is much simpler.
\end{rem}

\section{Proof of Theorem \ref{thm+three}}
\begin{proof}
The proof for the second part is similar to that of the first.
Therefore, for brevity, we only show that the linear transformation
$$y_n=\sum_{i=0}^{kn}\binom{n,k}{j}x_i$$ preserves the log-concavity
property.
We first prove the next fact.
\begin{f}\label{q+sum}
If $(x_n)_{n\geq0}$ is log-concave, then so is
$(\sum_{i=0}^kb_ix_{n+i})_{n\geq0}$.
\end{f}

\begin{proof}
Assume that $z_n= \sum_{i=0}^kb_ix_{n+i}$. Let
$$B=[b_{j-i}]_{i,j\geq0}=\left[\begin{array}{ccccccc}
b_0&b_{1}&b_{2}&\ldots&b_{k}&0&\ldots\\
0&b_0&b_{1}&\ldots&b_{k-1}&b_{k}&\ldots\\
0&0&b_0&\ldots&b_{k-2}&b_{k-1}&\ldots\\
\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ldots
\end{array}\right]$$
and $$X=[x_{j+i}]_{i,j\geq0}=\left[\begin{array}{ccccccc}
x_0&x_{1}&x_{2}&x_{3}&x_{4}&x_5&\ldots\\
x_{1}&x_{2}&x_{3}&x_{4}&x_5&\ldots&\ldots\\
x_{2}&x_{3}&x_{4}&x_5&\ldots&\ldots&\ldots\\
x_{3}&x_{4}&x_5&\ldots&\ldots&\ldots&\ldots\\
x_{4}&x_5&\ldots&\ldots&\ldots&\ldots&\ldots\\
\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ldots
\end{array}\right].$$
Let $Z=[z_{j+i}]_{i,j\geq0}$. It is clear that $Z=BX$. Note that $B$
is TP$_2$ since $(b_i)_{0\leq i\leq k}$ is log-concave and every
minor of order $2$ of $X$ is nonpositive because $(x_n)_{n\geq0}$ is
log-concave. So the classical Cauchy-Binet Theorem shows that every
minor of order $2$ of the product $Z$ of $B$ and $X$ is nonpositive.
Hence $(z_n)_{n\geq0}$ is log-concave. This completes the proof.
\end{proof}

In the following, we will prove that the linear transformation
$$y_n=\sum_{i=0}^{kn}\binom{n,k}{i}x_i$$ preserves the log-concavity
property by induction on $n$.

If $0\leq n \leq 2$, then by $y_n=\sum_{i=0}^{kn}\binom{n,k}{i}x_i$,
we obtain that
\begin{eqnarray}
y_1^2-y_2y_0
&=&\left[\sum_{i=0}^k\binom{1,k}{i}x_i\right]^2-x_0\sum_{i=0}^{2k}\binom{2,k}{i}x_i\nonumber\\
&=&\sum_{i=0}^{2k}\left[\sum_{j=0}^i\binom{1,k}{i-j}\binom{1,k}{j}x_jx_{i-j}\right]-x_0\sum_{i=0}^{2k}\sum_{j=0}^i\binom{1,k}{i-j}\binom{1,k}{j}x_i\nonumber\\
&\geq&\sum_{i=0}^{2k}\left[\sum_{j=0}^i\binom{1,k}{i-j}\binom{1,k}{j}x_0x_{i}\right]-\sum_{i=0}^{2k}\sum_{j=0}^i\binom{1,k}{i-j}\binom{1,k}{j}x_ix_0\nonumber\\
&\geq&0
\end{eqnarray}
because $x_jx_{i-j}\geq x_0x_i$ from the log-concavity of
$(x_n)_{n\geq0}$.

 Thus, we get that $y_0,y_1,y_2$ is log-concave. So we proceed to the
inductive step ($n\geq 3$).

Note that
\begin{eqnarray*}
y_n&=&\sum_{i=0}^{kn}\binom{n,k}{i}x_i\\
&=&\sum_{i=0}^{kn}\binom{n-1,k}{i}\sum_{j=0}^kb_jx_{i+j}.
\end{eqnarray*}
Thus by the induction hypothesis and the log-concavity of
$(\sum_{j=0}^kb_jx_{i+j})_{i\ge 0}$ by means of Fact \ref{q+sum}, we
have the sequence $(y_n)_{n\geq0}$ is log-concave. This completes
the proof.
\end{proof}
\section{Strong $q$-log-concavity of rows in combinatorial triangles}
In this section, we will give some sufficient conditions for the
strong $q$-log-concavity of rows in certain triangular arrays.
\begin{thm}\label{thm+q+log-concave}
 Assume that the infinite lower triangular array
$[A_{n,k}(q)]_{n,k\geq 0}$ satisfies the recurrence
\begin{eqnarray}\label{re}
A_{n,k}(q)=f_{n,k}(q)\,A_{n-1,k-1}(q)+g_{n,k}(q)\,A_{n-1,k}(q)+h_{n,k}(q)\,A_{n-1,k+1}(q)
\end{eqnarray}
for $n\geq 1$ and $k\geq 0$, where $A_{0,0}(q)=1$,
$A_{0,k}(q)=A_{0,-1}(q)=0$ for $k>0$. Assume that three sequences of
polynomials with nonnegative coefficients $(f_{n,k}(q))_{k\geq0}$,
$(g_{n,k}(q))_{k\geq0}$ and $(h_{n,k}(q))_{k\geq0}$ are strongly
$q$-log-concave in $k$, respectively. If
\begin{eqnarray*}
&&f_{n,l}(q)g_{n,k}(q)+f_{n,k}(q)g_{n,l}(q)\geq_q f_{n,l+1}(q)g_{n,k-1}(q)+f_{n,k-1}(q)g_{n,l+1}(q),\\
&&g_{n,l}(q)h_{n,k}(q)+g_{n,k}(q)h_{n,l}(q)\geq_q g_{n,l+1}(q)h_{n,k-1}(q)+g_{n,k-1}(q)h_{n,l+1}(q),\\
&&f_{n,l}(q)h_{n,k}(q)+f_{n,k}(q)h_{n,l}(q)\geq_qf_{n,k-1}(q)h_{n,l+1}(q)+f_{n,l+1}(q)h_{n,k-1}(q)\\
\text{and}&&g_{n,k}(q)g_{n,l}(q)\geq_q f_{n,l+1}(q)h_{n,k-1}(q)
\end{eqnarray*}
for all $1\leq k\leq l\leq n$, then, for any fixed $n$, the row
sequence $(A_{n,k}(q))_{0\leq k\leq n}$ is strongly $q$-log-concave
in $k$. In particular, each row sequence $(A_{n,k}(q))_{0\leq k\leq
n}$ is log-concave in $k$ for any fixed $q\geq0$.
\end{thm}

\begin{proof}
In order to prove that $(A_{n,k}(q))_{0\leq k\leq n}$ is strongly
$q$-log-concave in $k$, we only need to prove
$$A_{n,l}(q)A_{n,k}(q)- A_{n,l+1}(q)A_{n,k-1}(q)\geq_q0$$ for any $l\geq k$, which will be done by induction
on $n$. It is obvious for $n=0$. So, we assume that it follows for
$n\leq m-1$.  For brevity, we write $f_k$ (resp. $g_k$, $h_k$) for
$f_{n,k}(q)$ (resp. $g_{n,k}(q)$, $h_{n,k}(q)$). Then for $n=m$ and
$1\leq k\leq l\leq m$, by (\ref{re}), we have
\begin{eqnarray}\label{eq+main}
&&A_{m,l}(q)A_{m,k}(q)- A_{m,l+1}(q)A_{m,k-1}(q)\nonumber\\
&=&\left[f_{l}A_{m-1,l-1}(q)+g_{l}A_{m-1,l}(q)+h_{l}A_{m-1,l+1}(q)\right]\times\nonumber\\
&&\left[f_{k}A_{m-1,k-1}(q)+g_{k}A_{m-1,k}(q)+h_{k}A_{m-1,k+1}(q)\right]-\nonumber\\
&&\left[f_{l+1}A_{m-1,l}(q)+g_{l+1}A_{m-1,l+1}(q)+h_{l+1}A_{m-1,l+2}(q)\right]\times\nonumber\\
&&\left[f_{k-1}A_{m-1,k-2}(q)+g_{k-1}A_{m-1,k-1}(q)+h_{k-1}A_{m-1,k}(q)\right]\nonumber\\
&=&\underbrace{f_{l}f_{k}A_{m-1,l-1}(q)A_{m-1,k-1}(q)-f_{l+1}f_{k-1}A_{m-1,l}(q)A_{m-1,k-2}(q)}+\nonumber\\
&&f_{l}g_kA_{m-1,l-1}(q)A_{m-1,k}(q)-f_{k-1}g_{l+1}A_{m-1,l+1}(q)A_{m-1,k-2}(q)+\nonumber\\
&&[f_{k}g_{l}-f_{l+1}g_{k-1}]A_{m-1,l}(q)A_{m-1,k-1}(q)+\nonumber\\
&&\underbrace{h_{l}h_{k}A_{m-1,l+1}(q)A_{m-1,k+1}(q)-h_{l+1}h_{k-1}A_{m-1,l+2}(q)A_{m-1,k}(q)}+\nonumber\\
&&h_{k}g_{l}A_{m-1,l}(q)A_{m-1,k+1}(q)-h_{l+1}g_{k-1}A_{m-1,l+2}(q)A_{m-1,k-1}(q)+\nonumber\\
&&[h_{l}g_{k}-h_{k-1}g_{l+1}]A_{m-1,l+1}(q)A_{m-1,k}(q)+\nonumber\\
&&\underbrace{g_{l}g_{k}A_{m,l}(q)A_{m,k}(q)-g_{l+1}g_{k-1}A_{m-1,l+1}(q)A_{m-1,k-1}(q)}+\nonumber\\
&&f_{l}h_{k}A_{m-1,l-1}(q)A_{m-1,k+1}(q)-h_{l+1}f_{k-1}A_{m-1,l+2}(q)A_{m-1,k-2}(q)+\nonumber\\
&&[h_{l}f_{k}A_{m-1,l+1}(q)A_{m-1,k-1}(q)-f_{l+1}h_{k-1}A_{m-1,l}(q)A_{m-1,k}(q)].
\end{eqnarray}
In what follows we will prove the nonnegativity of (\ref{eq+main})
in $q$.

Firstly, it follows from the strong $q$-log-concavities of
$(A_{m-1,k}(q))_{0\leq k\leq m-1}$ and $(f_k(q))_{k\geq0}$ that
 \begin{eqnarray}\label{ieq1}A_{m-1,l-1}(q)A_{m-1,k-1}(q)-A_{m-1,l}(q)A_{m-1,k-2}(q)\geq_q0\nonumber\end{eqnarray}
 and \begin{eqnarray}\label{ieq1}f_{l}f_{k}-f_{l+1}f_{k-1}\geq_q0,\nonumber\end{eqnarray}
which implies
\begin{eqnarray}\label{ieq1}f_{l}f_{k}A_{m-1,l-1}(q)A_{m-1,k-1}(q)-f_{l+1}f_{k-1}A_{m-1,l}(q)A_{m-1,k-2}(q)\geq_q0.\end{eqnarray}
Similarly, we also have
\begin{eqnarray}\label{ieq2}h_{l}h_{k}A_{m-1,l+1}(q)A_{m-1,k+1}(q)-h_{l+1}h_{k-1}A_{m-1,l+2}(q)A_{m-1,k}(q)\geq_q0.\end{eqnarray}

Secondly, by the strong $q$-log-concavity of $(A_{m-1,k}(q))_{0\leq
k\leq m-1}$ and $$f_{l}g_{k}+f_{k}g_{l}\geq_q
f_{l+1}g_{k-1}+f_{k-1}g_{l+1},$$ we get
\begin{eqnarray}\label{ieq4}
&&f_{l}g_kA_{m-1,l-1}(q)A_{m-1,k}(q)-f_{k-1}g_{l+1}A_{m-1,l+1}(q)A_{m-1,k-2}(q)+\nonumber\\
&&[f_{k}g_{l}-f_{l+1}g_{k-1}]A_{m-1,l}(q)A_{m-1,k-1}(q)\nonumber\\
&\geq_q&[f_{l}g_k+f_{k}g_{l}-f_{l+1}g_{k-1}]A_{m-1,l}(q)A_{m-1,k-1}(q)-f_{k-1}g_{l+1}A_{m-1,l+1}(q)A_{m-1,k-2}(q)\nonumber\\
&\geq_q&[f_{l}g_k+f_{k}g_{l}-f_{l+1}g_{k-1}-f_{k-1}g_{l+1}]A_{m-1,l}(q)A_{m-1,k-1}(q)\nonumber\\
&\geq_q&0.
\end{eqnarray}
In a similar way, we also have
\begin{eqnarray}\label{ieq5}
&&h_{k}g_{l}A_{m-1,l}(q)A_{m-1,k+1}(q)-h_{l+1}g_{k-1}A_{m-1,l+2}(q)A_{m-1,k-1}(q)+\nonumber\\
&&[h_{l}g_{k}-h_{k-1}g_{l+1}]A_{m-1,l+1}(q)A_{m-1,k}(q)\nonumber\\
&\geq_q&[h_{k}g_{l}+h_{l}g_{k}-h_{k-1}g_{l+1}-h_{l+1}g_{k-1}]A_{m-1,l+1}(q)A_{m-1,k}(q)\nonumber\\
&\geq_q&0.
\end{eqnarray}
Finally,  it follows from the strong $q$-log-concavity of
$(g_k(q))_{k\geq0}$ that
\begin{eqnarray}\label{ieq1}g_{l}g_{k}-g_{l+1}g_{k-1}\geq_q0.\nonumber\end{eqnarray}
Thus we have
\begin{eqnarray}\label{ieq6}
&&\underbrace{g_{l}g_{k}A_{m-1,l}(q)A_{m-1,k}(q)-g_{l+1}g_{k-1}A_{m-1,l+1}(q)A_{m-1,k-1}(q)}+\nonumber\\
&&f_{l}h_{k}A_{m-1,l-1}(q)A_{m-1,k+1}(q)-h_{l+1}f_{k-1}A_{m-1,l+2}(q)A_{m-1,k-2}(q)+\nonumber\\
&&[h_{l}f_{k}A_{m-1,l+1}(q)A_{m-1,k-1}(q)-f_{l+1}h_{k-1}A_{m-1,l}(q)A_{m-1,k}(q)]\nonumber\\
&\geq_q&\underbrace{g_{l}g_{k}[A_{m-1,l}(q)A_{m-1,k}(q)-A_{m-1,l+1}(q)A_{m-1,k-1}(q)]}+\nonumber\\
&&f_{l}h_{k}A_{m-1,l-1}(q)A_{m-1,k+1}(q)-h_{l+1}f_{k-1}A_{m-1,l+2}(q)A_{m-1,k-2}(q)+\nonumber\\
&&\underbrace{[f_{l+1}h_{k-1}+h_{l+1}f_{k-1}-f_{l}h_{k}]}A_{m-1,l+1}(q)A_{m-1,k-1}(q)-\nonumber\\
&&f_{l+1}h_{k-1}A_{m-1,l}(q)A_{m-1,k}(q)]\nonumber\\
&=&\underbrace{g_{l}g_{k}[A_{m-1,l}(q)A_{m-1,k}(q)-A_{m-1,l+1}(q)A_{m-1,k-1}(q)]}+\nonumber\\
&&\underbrace{f_{l+1}h_{k-1}[A_{m-1,l+1}(q)A_{m-1,k-1}(q)-A_{m-1,l}(q)A_{m-1,k}(q)]}+\nonumber\\
&&f_{l}h_{k}[A_{m-1,l-1}(q)A_{m-1,k+1}(q)-A_{m-1,l+1}(q)A_{m-1,k-1}(q)]+\nonumber\\
&&\underbrace{h_{l+1}f_{k-1}[A_{m-1,l+1}(q)A_{m-1,k-1}(q)-A_{m-1,l+2}(q)A_{m-1,k-2}(q)]}\nonumber\\
&\geq_q&\underbrace{[g_{l}g_{k}-f_{l+1}h_{k-1}][A_{m-1,l}(q)A_{m-1,k}(q)-A_{m-1,l+1}(q)A_{m-1,k-1}(q)]}\nonumber\\
&\geq_q&0
\end{eqnarray}
since
$$f_{l}h_{k}+f_{k}h_{l}\geq_qf_{k-1}h_{l+1}+f_{l+1}h_{k-1}$$
and $$g_kg_l\geq_q f_{l+1}h_{k-1}.$$

Thus, by (\ref{eq+main})- (\ref{ieq6}), we get
$$A_{m,k}(q)A_{m,l}(q)- A_{m,l+1}(q)A_{m,k-1}(q)\geq_q 0$$
for $1\leq k\leq l\leq m$. The proof is complete.
\end{proof}

The following special case related to the Riordan array
\cite{MRSV97} may be more interesting.
\begin{prop}
Define the matrix $[A_{n,k}(q)]_{n,k\geq0}$ recursively:
\begin{eqnarray*}
A_{0,0}(q)&=&1, \quad A_{0,k}(q)=0\quad(k>0),\\
A_{n,0}(q)&=&e(q)\,A_{n-1,0}(q)+h(q)\,A_{n-1,1}(q),\\
A_{n,k}(q)&=&A_{n-1,k-1}(q)+g(q)\,A_{n-1,k}(q)+h(q)\,A_{n-1,k+1}(q)\quad
(n,k\geq 1).
\end{eqnarray*}
If $e(q)g(q)\geq_q h(q) \geq_q 0$ and $g(q)\geq_q e(q)\geq_q 0$,
then each row $(A_{n,k}(q))_{0\leq k\leq n}$ is strongly
$q$-log-concave.
\end{prop}

Applying Theorem \ref{thm+q+log-concave} to combinatorial arrays in
Example \ref{exm}, we have the following results in a unified
manner.
\begin{cor}
Each row sequence of the Bell polynomials triangle, the Eulerian
polynomials triangle and the Narayana polynomials triangle is
strongly $q$-log-concave, respectively.
\end{cor}
\begin{cor}
Each row sequence in the Catalan triangles of Aigner and Shaprio,
the Motzkin triangle, the large Schr\"oder triangle, and the Bell
triangle is log-concave, respectively.
\end{cor}


\end{document}